OpenStudy (anonymous):

I have to find the range of the height of a pine tree. its standing up beside a right triangle. The shadow of the tree is 8 ft. The hypotenuse is 17 ft. How would I do this to figure out my answer?

OpenStudy (anonymous):

Pythagorean theorem a²+b²=c²

OpenStudy (anonymous):

\( \large 8^{2} + x^{2}=17^{2}\)

OpenStudy (anonymous):

So how would I find the range of the pine tree?

OpenStudy (anonymous):

if you find the value of x, you'll know the height of the tree

OpenStudy (anonymous):

So i fid x by subtracting 8 from both sides?

OpenStudy (anonymous):

that would be a good start

OpenStudy (anonymous):

okay so I did 8^2-8^2 which is 0. so that cancels out so al i have left is x^2=225. because I did 17^2-8^2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so if \[\large x^{2}=225 \] then x=?

OpenStudy (anonymous):

x=225?

OpenStudy (anonymous):

if you go from x² to x, you do \[\large \sqrt{x²} = x\]

OpenStudy (anonymous):

so if you take the root of one side of the = sign, you have to do it on the other side as well

OpenStudy (anonymous):

x=15

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

So how would I find the range of the height of the triangle. I have to do the triangle inequality theorem.

OpenStudy (anonymous):

Eh, i don't understand the question

OpenStudy (anonymous):

It says find the range of the pine tree by using the triangle inequality theorem.

OpenStudy (anonymous):

I think i misinterpret the question then, hold on.

OpenStudy (anonymous):



OpenStudy (anonymous):

yes but that's not really the triangle inequality theorem, i never used that before either.

OpenStudy (anonymous):

Well, it says to find the range of the pine tree. but i can refer to the triangle inequality theroem.

OpenStudy (anonymous):

Ok, so i googled a around a bit and apperantly x+8>18 so the tree can't be smaller then 10, but we don't know the upper limit yet then... http://www.mathopenref.com/triangleinequality.html

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

oh wait it;'s obvious actually

OpenStudy (anonymous):

it can never be longer then 18+8 either i guess, so it's between 10 and 26 i guess

OpenStudy (anonymous):

at least f i understood the theorem correctly.

OpenStudy (anonymous):

Thats good. It really helped me. I need to find all 3 triginometric ratios using angle a.